OpenStudy (anonymous):

logbasea=507.2 and logbaseb=505.2 what is a/b?

OpenStudy (amistre64):

your missing arguements

OpenStudy (amistre64):

\[log_a=507.2 \] the notationis a bit odd to me

OpenStudy (anonymous):

my bad, there is no given base

OpenStudy (anonymous):

is 10

OpenStudy (anonymous):

I mean the base

OpenStudy (amistre64):

\[L(a)=M;\ L(b)=(N)\] \[a=10^M;\ b=10^N\] \[\frac ab=\frac{10^M}{10^N}=10^{M-N}\]

OpenStudy (amistre64):

usually log represents base 10; but in some cases it means base "e" so these days you have to be careful

OpenStudy (amistre64):

if it aint \(10^{M-N}\) then it should be \(e^{M-N}\) but im pretty confident its the 10 spot :)